Michael, a hamster lover, keeps his four pet hamster's in a roomy cage, where they spend much of their spare time–when they're not sleeping or eating–joyfully scampering about on the cage's floor. Michael tracks his hamster's health diligently and just now recorded their masses as 0.0211 kg, 0.0163 kg, 0.0247 kg, and 0.0167 kg. At this very instant the x and y components of the hamster's velocities are, respectively, (0.957 m/s, -0.513 m/s), (-0.827 m/s, -0.873 m/s), (0.233 m/s, 0.311 m/s), and (-0.101 m/s, 0.427 m/s). Calculate the x and y components of Michael's hamster's total momentum, px and py.

px = ____ kg m/s

py = ____ kg m/s
Answer #1

m1= 0.0211 kg, m2= 0.0163 kg, m3 =0.0247 kg, m4= 0.0167 kg

v1= (0.957 m/s, -0.513 m/s), v2= (-0.827 m/s, -0.873 m/s), v3= (0.233 m/s, 0.311 m/s), v4=(-0.101 m/s, 0.427 m/s)

Px = P1x + P2x +P3x +P4x = m1v1x + m2v2x + m3v3x + m4v4x

Px = (0.0211kg*0.957m/s) + (0.0163 kg*-0.827m/s) +(0.0247kg*0.233 m/s) +(0.0167kg*-0.101 m/s)

Px = 0.0108 kg*m/s

Py = P1y + P2y +P3y +P4y = m1v1y + m2v2y + m3v3y + m4v4y

Py = (0.0211kg*-0.513 m/s) + (0.0163 kg*-0.873m/s) +(0.0247kg*0.311 m/s) +(0.0167kg*0.427 m/s)

Py = -0.0103 kg*m/s